A line passes through (–9, –5) and (2, –4).
a. Write an equation for the line in point-slope form.
b. Rewrite the equation in standard form using integers.

Answer by TimothyLamb(4379)

first use the two-point form:
(–9, –5) and (2, –4)
x1 y1 x2 y2 = –9 –5 2 –4
plug x1 y1 x2 y2 into: https://sooeet.com/math/linear-equation-solver.php
---
slope (m) = 0.0909090909 ~= 0.091 = 91/1000
y-intercept = -4.18181818 ~= -4.181
---
write the equation in point-slope form:
y + 5 = 0.091(x + 9)
---
rewrite the equation in standard form using integers:
y + 5 = (91/1000)(x + 9)
y + 5 = (91/1000)x + 819/1000
(91/1000)x + (-1)y = 5 - 819/1000
(91/1000)x + (-1)y = 5000/1000 - 819/1000
(91/1000)x + (-1)y = 4181/1000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php